# Global Hash Tables Strike Back! An Analysis of Parallel GROUP BY Aggregation

### v1

@@ Abstract.
Efficiently computing group aggregations (i.e., GROUP BY) on modern many-core architectures is critical for analytic database systems. Today’s engines predominately use a partitioned approach to group aggregation, in which an incoming data stream is partitioned by key values so that every row for a particular key is sent to the same thread. In this paper, we revisit a simpler strategy: a fully concurrent group aggregation technique using a shared global hash table. While approaches using general-purpose concurrent hash tables have generally been found to perform worse than partitioning-based approaches, we argue that the key ingredient is customizing the concurrent hash table for the specific task of group aggregation. Through extensive experiments on synthetic workloads (varying key cardinality, skew, and thread counts), we demonstrate that a purpose-built concurrent hash table can match or surpass partitioning-based techniques. We also analyze the operational characteristics of both techniques, including resizing costs and memory pressure. In the process, we derive practical guidelines for database implementers. Overall, our analysis indicates that fully concurrent group aggregation is a viable alternative to partitioning.
Artifact Availability:
The source code, data, and/or other artifacts have been made available at https://github.com/danielxue/global-hash-tables-strike-back.
@@ 1. Introduction
Group aggregation, represented in SQL via GROUP BY, is a fundamental operation in analytical query processing, especially decision-support workloads (Nambiar and Poess, 2006). To ensure that database systems continue to scale well with new many-core architectures, it is critical to build highly concurrent group aggregation schemes.
Analytic database systems today are quite diverse. For example, Datafusion (Lamb et al., 2023) follows a Volcano-style (Graefe and McKenna, 1993) block iteration approach (i.e., “pull”), where as DuckDB (Raasveldt and Mühleisen, 2019) follows a HyPer-inspired morsel driven parallelism (Leis et al., 2014) approach (i.e., “push”). Despite drastic differences in their execution models, nearly all of today’s analytic database systems use partitioning techniques in order to parallelize group aggregations. To the best of our knowledge, the partitioning approach proposed by Leis et al. (Leis et al., 2014), which combines local preaggregation with partitioning (we provide background on partitioning-based techniques in Section 2.2), has become dominant in many modern analytic systems (Kuiper et al., 2024; Lamb et al., 2023).
@@
In this paper, we explore the design space of group aggregation algorithms using a global concurrent hash table, comparing against state-of-the-art partitioning approaches. Most significantly, we find that a simple, purpose-built concurrent hash table using linear probing and a customized get-or-insert function can not only scale well on modern multi-core hardware but can match or even outperform partitioning-based approaches. Our results do not show that one approach is better than the other, but instead highlight the operational benefits and costs of both approaches. Throughout our exploration, we make specific recommendations for database implementers.
Our implementation of group aggregation with a global concurrent hash table closely tracks the implementation in MonetDB (Boncz et al., 2005). Each worker, upon receiving a row, first obtains from a hash table an integer “ticket” for that row’s grouping key. This ticket uniquely identifies each group and serves as an index to locate the aggregated value for that group. We provide a detailed description of this procedure in Section 2.3. This two-phase procedure: ticketing and partial aggregate update, is repeated for each row (possibly in a vectorized fashion: ticketing an entire morsel, then aggregating that morel), and opens up a large number of possible designs, which we explore.
@@
In the initial phase of fully concurrent group aggregation, each unique group is assigned an integer “ticket.” This operation can be performed with a concurrent hash table that atomically: checks if a key is already in the table, returning the ticket for that key if so, and, if not, inserting a new ticket into the table. Surprisingly, many general-purpose hash tables cannot perform this operation atomically or do not optimize for this particular case. As a result, there is significant room for improvement by building a specialized fast path for this particular operation, while avoiding adding extra overhead for unnecessary operations like deletes. We test several implementations based on atomics and fine-grained locking, and show that simple purpose-built hash tables can significantly outperform their complex general-purpose counterparts.
@@
After ticketing, each worker must apply the relevant aggregation function on the values associated with each row. This can be done in either a thread-specific way (e.g., each thread maintains local aggregation storage, merging the results at the end), or using concurrent access to global space. We explore the tradeoffs between both approaches, characterizing their operational tradeoffs. Specifically, we find that concurrent access to global space works well in the absence of heavy hitters, and we propose a simple thread-local approach that works well except for when every grouping key is unique.
@@
Our experimental study is organized as follows. In Section 2, we explain our assumed model of execution, and introduce the basics of partitioned aggregation and concurrent aggregation. In Section 3, we investigate the design space of aggregation with a concurrent hash table. In Section 3.1, we explore the ticketing step, and in Section 3.2, we explore the aggregation step. In Section 4, we investigate the end-to-end performance of both approaches. We analyze their scaling properties in a number of different scenarios, including a causal analysis of scaling behavior (i.e., identifying bottlenecks). We additionally discuss several important operational characteristics, like memory usage and hash table resizing. In Section 5 we discuss related works, before concluding and outlining future work in Section 6.
@@ 2. Preliminaries
In this section, we first establish the model of query execution we are operating within and discuss the relevant constraints it places upon our work. We then provide an overview of the two models of aggregation that we assess in this paper: partitioned and fully concurrent.
@@ 2.1. Model of Execution
Leis et al. (Leis et al., 2014) introduced the morsel-driven framework of parallel query execution to improve performance in main-memory systems where latency is compute-bound rather than I/O-bound. In this model, horizontal parallelism is achieved by breaking down work into data fragments called “morsels.” Morsels are dynamically assigned to distributed to threads in a pool, using strategies such as work-stealing to ensure even distribution. In this model, query execution is pipelined, with morsels being pushed to the next operator as they are finished being processed. Morsel-driven parallelism is also often used with columnar data representations (Leis et al. integrated their work within HyPer and used its columnar data representation), with morsels serving as units for vector-at-a-time execution. Operating on a dense vector of values enables important optimizations such as amortizing interpretation overhead (e.g. dynamic type resolution) and SIMD instructions (Kersten et al., 2018).
The morsel-driven model places some restrictions on operator implementations. Operators must be able to operate on chunks of data at a time to conform with pipelining and do not have access to the rest of the incoming tuples. This restriction is especially problematic for sort-based aggregations, although some variations propose ways to adapt to such models of execution (Müller et al., 2015). Due to the prevalence of main-memory and morsel-driven systems, we focus our investigation on only operator implementations that conform to this model of execution.
@@ 2.2. Partitioned Aggregation Model
Under a partitioned approach to parallelized aggregation, concurrency control is avoided by assigning a subset of the key domain to each thread. A naive partitioning strategy assigns incoming tuples to their proper threads, which then aggregates the assigned tuples locally. However, this method suffers significantly from data skew and has been found to perform worse than methods utilizing a local hash table to preaggregate heavy hitters before partitioning (Ye et al., 2011).
Figure 1. A sample execution of partitioned aggregation. This diagram is modified slightly from the one presented in (Leis et al., 2014). This example is given for demonstrative purposes and to contrast with the sample execution on the same input in Figure 2 of our fully concurrent model.
Leis et al. (Leis et al., 2014) describe a specific partitioned aggregation strategy illustrated in Figure 1 to parallelize group by aggregations within morsel-driven systems. Their method consists of two stages: local preaggregation and partition-wise aggregation In the first stage, each thread aggregates all values from their assigned morsel(s) in a local, fixed-sized hash table ❶. When the hash tables become full, the partially aggregated values (referred to as partial aggregates) are spilled into partitions ❷ and the process continues. After all data is preaggregated, all partial aggregates are flushed to their proper partition. In the second stage, partitions are exchanged between threads and each thread now aggregates all the partial aggregates from all other threads to compute the final answer ❸.
This algorithm is motivated by the need for skew resistance in partitioned aggregation methods, avoiding uneven work distribution by spreading the work of aggregating heavy hitters among all threads in the local preaggregation stage. However, for high cardinality workloads, there is repeated spilling from the local aggregation tables. The constant spilling results in each tuple essentially being aggregated twice, once in each stage of the algorithm, creating a significant source of overhead.
The local preaggregation approach to partitioned aggregation has been adopted by a number of real-world systems including DuckDB (Kuiper et al., 2024) and Datafusion (Lamb et al., 2023). Due to the widespread adoption of this particular algorithm and its known good scaling behavior (Ye et al., 2011; Kuiper et al., 2024), it is used as the baseline partitioned aggregation method against which we compare the performance of our fully concurrent aggregation algorithm.
@@ 2.3. Fully Concurrent Aggregation Model
Figure 2. A sample execution of our fully concurrent aggregation model using the same instance as in Figure 1. This particular diagram aligns most closely to our atomic or locking method of partial aggregate updates. In contrast to partitioned aggregation, instead of completing all work in local data structures, there are now two shared structures, the ticketing hash table and the vector of partial aggregates.
To perform a group by aggregation in a fully concurrent manner, each thread must aggregate an arbitrary morsel of data from start to finish with any key distribution. We separate fully concurrent aggregation into two different steps: ticketing and partial aggregation update (which we will often refer to as just the “update” step), as shown in Figure 2.
In the ticketing step ❶, a concurrent shared hash table is used to map each key value to an integer “ticket.” This mapping is one-to-one: each unique key is granted a single, unique ticket, and the ticket assigned to a key is consistent across all threads due to the use of a shared hash table. A ticket conceptually represents the location where the corresponding partial aggregate to be updated is stored. In our implementation, this is an index into a vector of partial aggregates that are later updated based on the value being aggregated. Another view of the ticketing step is that it uses a hash table to incrementally create a perfect hash function over the key space. This stage forms the “group by” part of group by aggregations.
Next, the resulting vector of tickets and the value column is fed to the update step ❷ where the query’s specified aggregation function(s) are applied to update each partial aggregate (e.g. incremented for COUNT, added to for SUM, etc). It is easiest to view partial aggregates as a global vector with cells protected individually with some concurrency-safe data structure (a lock or atomic), although we do later introduce a thread local update procedure that does not align with this conceptualization. This step is the “aggregation” part of group by aggregation.
@@
Note that this layer of indirection separating the partial aggregate from the hash map has already been common in practice for a myriad of reasons. For example, it is used by MonetDB  (Boncz et al., 2005) to enable vectorized execution, as well as by DuckDB (Kuiper et al., 2024) and DataFusion (Lamb et al., 2023). Indirection also enables optimizations only possible when acting on a dense column of values, such as the use of SIMD instructions or to amortize the cost of dynamic type resolution in non-compiled systems.
Given that indirection is already commonplace, we formalize it in our model to take advantage of its logical properties. In particular, since each unique key is assigned a ticket exactly once, our hash table workload is reduced to only lookups and inserts. One key contribution of this paper is noting that while concurrent hash tables have traditionally been eschewed in the context of aggregations due to poor scaling behavior with certain workloads, once constrained to inserts and lookups only, concurrent hash tables can be designed with excellent performance. This indirection pushes concurrent updates to the second stage. Crucially, concurrent updates now act on a vector of values rather than units of a hash table, allowing for alternate ways of protecting against simultaneous updates that do not degrade the performance of the ticketing hash table.
@@ 3. Fully Concurrent Aggregation
In this section, we investigate the design space for each stage of fully concurrent aggregation. We discuss various methods and provide micro-benchmarks of each stage in isolation to identify the best-performing method(s) for each stage, which we later test end-to-end in Section 4. We also discuss directions for future work to investigate that could leverage our particular model of aggregation to achieve more efficient operation compared to some of the existing, more general-purpose methods that we test.
@@ 3.1. Ticketing
When a new key arrives, an insert operation is required. Following a successful insertion for a key, all future requests should look up the previously returned ticket value. In many concurrent hash table designs, this lookup operation can be achieved with significantly less overhead than an insert (since shared locks are cheaper than exclusive locks, and simultaneous atomic reads are much cheaper than simultaneous atomic writes). However, since we do not know ahead of time whether an insert is necessary (i.e., we do not know if a particular key has already been given a ticket), it is crucial to provide an efficient fast path for lookup before attempting a more expensive insert. In doing so, we can greatly reduce contention effects from simultaneous requests on the same key, even when given a heavily skewed key distribution.
Some care in the implementation is needed to ensure correctness since during the gap between the fast path for lookup and the insertion attempt, another thread could have inserted the same key. A double insertion would violate the constraint that ticketing is bijective. We propose that the lookup and insert paths be integrated into one operation, which we denote GET_OR_INSERT, to ensure efficient and correct implementation. Such an operation is rarely supported out-of-the-box by concurrent hash maps, but depending on the specific hash map implementation the same effect could be achieved using an entry API or with a lookup followed by a non-overwriting insert.
@@
To efficiently retrieve keys in ticketed order when materializing our final results, we also need the hash table to store a copy of keys in ticket order. This creates some performance overhead but is generally efficient since there is no contention – only the thread issuing a ticket is responsible for storing a copy. The ticket-order copy of keys can also be maintained as the only copy to decrease memory overhead (e.g., since the ticket value can be used to lookup the key during hash map operations), but would come at the cost of an additional cache miss per lookup.
Generating Ticket Values. Generating a ticket value, while simple at face value, is a surprisingly non-trivial task due to synchronization concerns. Multiple threads must avoid issuing the same ticket to different keys. A naive implementation of a “ticketer” would use an atomic counter updated with a FETCH_ADD instruction whenever a thread needs a new ticket issued (on each insert). However, this introduces a significant source of contention for insert-heavy workloads. To combat this issue, one can use a fuzzy ticketer that assigns each thread a range of tickets to issue at a time. Each thread only needs to concurrently access an atomic value when it exhausts its assigned range, whose size can be tuned to all-but-eliminate contention. The tradeoff is that the vector of partial aggregates may no longer be perfectly dense; however, the number of gaps is bounded linearly by the number of threads, and, in most workloads, will be concentrated at the end of the aggregator vector, making them easy and efficient to eliminate at the end of the aggregation process.
Using a fuzzy ticketer also provides greater flexibility when implementing the insert operation. In cases where a ticket is obtained optimistically (e.g. as a parameter to an insert operation that fails when the entry already exists), a simple FETCH_ADD is insufficient because there is no guarantee the obtained ticket will actually be inserted, yet the underlying atomic counter has already been updated. For example, in our procedure in Algorithm 1, a ticket value is directly written to the table using a COMPARE_AND_SWAP (CAS) that may not succeed.
Figure 3. Performance of the Folklore* hash table when using a fuzzy ticketer as opposed to an incrementing atomic counter to issue tickets. Performance is evaluated with both low and high cardinality workloads and is measured with latency (lower is better).
There is a significant gap in latency on insert-heavy workloads between using a pure atomic counter as the ticketer and our fuzzy ticketer, as seen in Figure 3 (lower latency is better). This micro-benchmark is tested on our best-performing hash table design, Folklore*. Latency in the high cardinality workload is greater for the pure atomic method by a factor of 2.5x. The performance difference is negligible in the low cardinality workload (since there are few inserts in the first place). Note that the high cardinality workload here is not even the extreme case of being insert-only, but is merely $\sim$10insertion. Since using a single atomic ticket value drastically degrades performance when there are many unique keys, we recommend that implementers avoid this contention by using a structure that amortizes the cost of concurrent accesses across multiple inserts, such as our fuzzy ticketer.
Hash Table Designs
We benchmark a variety of state-of-the-art hash table designs. These include cuckoo hashing (Li et al., 2014), Iceberg (Pandey et al., 2023), Rust’s Leapfrog [^1] https://github.com/robclu/leapfrog library which uses leapfrog probing (Preshing, 2016) (which in turn is a variation on iceberg probing (Herlihy et al., 2008; Kelly et al., 2020)), and Rust’s popular DashMap. [^2] https://github.com/xacrimon/dashmap Our implementations additionally benefit from vectorized execution that optimizes hashing and amortizes the acquisition of a read lock on the shared table (to ensure the correctness of resizing).
We also implement a variant of the Folklore hash table proposed by Maier et al. (Maier et al., 2018), a lockless linear probing hash table. Our implementation leverages the lookup and insert-only workload and notes that ticket value 0 can be easily reserved as a sentinel “empty value”, which we relate to a corresponding “empty key.” These properties allow our design to use only a single word CAS instruction instead of the two-word version required by the canonical implementation (which is not a universally supported operation). We denote our implementation Folklore* to indicate this difference. The GET_OR_INSERT procedure for this design is outlined in Algorithm 1.
@@
     loop
         $(t,k):=(table[idx].t.load(),table.[idx].k.load())$
         if $t\neq 0$ then $\triangleright$ Fast path lookup.
@@
                  return $table[idx].t.load()$
              else if $k=EmptyKey$ then
                  continue
@@
         else if $table[idx].t.CAS(0,ticket)$ then $\triangleright$ Insert.
@@
              return ticket
@@
Evaluation

### v2

@@ Abstract.
Efficiently computing group aggregations (i.e., GROUP BY) on modern architectures is critical for analytic database systems. Hash-based approaches in today’s engines predominantly use a partitioned approach, in which incoming data is partitioned by key values so that every row for a particular key is sent to the same thread. In this paper, we revisit a simpler strategy: a fully concurrent aggregation technique using a shared hash table. While approaches using general-purpose concurrent hash tables have generally been found to perform worse than partitioning-based approaches, we argue that the key ingredient is customizing the concurrent hash table for the specific task of group aggregation. Through experiments on synthetic workloads (varying key cardinality, skew, and thread count), we demonstrate that in morsel-driven systems, a purpose-built concurrent hash table can match or surpass partitioning-based techniques. We also analyze the operational characteristics of both techniques, including resizing costs and memory pressure. In the process, we derive practical guidelines for database implementers. Overall, our analysis indicates that fully concurrent group aggregation is a viable alternative to partitioning.
PVLDB Artifact Availability:
The source code, data, and/or other artifacts have been made available at %leave␣empty␣if␣no␣availability␣url␣should␣be␣sethttps://github.com/danielxue/global-hash-tables-strike-back.
@@ 1. Introduction
Group aggregation, represented in SQL via GROUP BY, is a fundamental operation in analytical query processing, especially decision-support workloads (Nambiar and Poess, 2006). To ensure that database systems continue to scale well with new many-core architectures, it is critical to build highly concurrent group aggregation schemes. While many different schemes for group aggregation exist (e.g., sorting), this paper will analyze hash-based aggregation techniques.
Analytic database systems today are quite diverse. For example, Datafusion (Lamb et al., 2023) follows a Volcano-style (Graefe and McKenna, 1993) block iteration approach (i.e., “pull”), whereas DuckDB (Raasveldt and Mühleisen, 2019) follows a HyPer-inspired morsel-driven parallelism (Leis et al., 2014) approach (i.e., “push”). Despite drastic differences in their execution models, nearly all of today’s analytic database systems use partitioning techniques to parallelize group aggregations. To the best of our knowledge, the partitioning approach proposed by Raman et al.  (Raman et al., 2013) has become dominant in many modern analytic systems (Leis et al., 2014; Kuiper et al., 2024; Lamb et al., 2023), which uses preaggregation to alleviate pressure from data skew. We provide background on partitioning-based techniques in Section 2.2.
@@
In this paper, we explore the design space of group aggregation algorithms using a global concurrent hash table, comparing against the state-of-the-art partitioned approach in morsel-driven (Leis et al., 2014) execution engines. Most significantly, we find that a simple, purpose-built concurrent hash table using linear probing and a customized get-or-insert function can scale well on modern multi-core hardware and match or even outperform partitioning-based approaches. Our results do not show that one approach is better than the other, but instead highlight the operational benefits and costs of both approaches. Throughout our exploration, we make specific recommendations for database implementers.
Our implementation of group aggregation with a global concurrent hash table closely tracks the implementation in MonetDB (Boncz et al., 2005). Each worker, upon receiving a row, first obtains from a hash table an integer “ticket” for that row’s grouping key. This ticket uniquely identifies each group and serves as an index to locate the aggregated value for that group. We provide a detailed description of this procedure in  Section 2.3. This two-phase procedure, ticketing and partial aggregate update, is repeated for each row, and opens up a large number of possible designs, which we explore.
@@
In the initial phase of fully concurrent group aggregation, each unique group is assigned an integer “ticket.” This operation can be performed with a concurrent hash table that atomically checks if a key is already in the table, returning the ticket for that key if so, and, if not, inserting a new ticket into the table. Surprisingly, many general-purpose hash tables cannot perform this operation atomically or do not optimize for this particular case. As a result, there is significant room for improvement by building a specialized fast path for this particular operation, while avoiding adding extra overhead for unnecessary operations like deletes. We test several implementations based on atomics and fine-grained locking, and show that simple purpose-built hash tables can significantly outperform their complex general-purpose counterparts.
@@
After ticketing, each worker must apply the relevant aggregation function on the values associated with each row. This can be done in either a thread-specific way (e.g., each thread maintains local aggregation storage, merging the results at the end), or using concurrent access to global space. We explore the tradeoffs between both approaches, characterizing their operational tradeoffs. Specifically, we find that concurrent access to global space works well in the absence of heavy hitters, and we propose a simple thread local approach that works well except for when every grouping key is unique.
@@
Our experimental study is organized as follows. In Section 2, we explain our assumed model of execution, and introduce the basics of partitioned aggregation and concurrent aggregation. We also introduce the experimental setup used throughout our experiments. In Section 3, we investigate the design space of aggregation with a concurrent hash table. In Section 3.1, we explore the ticketing step, and in Section 3.2, we explore the aggregation step. In Section 4, we analyze the end-to-end performance of both approaches. In Section 5 we discuss related works, before concluding and outlining future work in Section 6.
@@ 2. Preliminaries
In this section, we first establish the model of query execution we are operating within and discuss the relevant constraints it places upon our work. We then provide an overview of the two models of aggregation assesed in this paper: partitioned and fully concurrent.
@@ 2.1. Model of Execution
Leis et al. (Leis et al., 2014) introduced the morsel-driven framework for parallel execution to improve performance in main-memory systems, where latency is compute-bound rather than I/O-bound. In this model, parallelism is achieved by breaking down work into data fragments called “morsels.” Morsels are dynamically assigned to threads in a pool, using work-stealing to ensure even distribution. A thread carries a morsel through the entire pipeline before grabbing another. Morsel-driven parallelism is often used with columnar data representations (e.g., HyPer (Leis et al., 2014)), with morsels serving as units for vector-at-a-time execution. Operating on a dense vector of values enables important optimizations, such as amortizing interpretation overhead and SIMD instructions (Kersten et al., 2018). However, the morsel-driven model places some restrictions on operator implementations. Operators must be able to operate on chunks of data at a time to conform with pipelining and cannot access the rest of the incoming tuples.
In this work, we focus on the morsel-driven execution model due to its popularity in open source OLAP systems (e.g., (Raasveldt and Mühleisen, 2019)). While our conclusions and recommendations may apply to other execution models (e.g., Datafusion (Lamb et al., 2023) uses operator-level parallelism but adopts an aggregation strategy from a morsel-driven system), different execution models may also lead to significant changes in performance properties. We leave investigations under alternative execution models to future work.
@@ 2.2. Partitioned Aggregation Model
In a partitioned approach, synchronization is avoided by splitting the key domain among threads. A naive partitioning strategy assigns incoming tuples to their proper thread (typically by taking the radix or mod of the key’s hash), which then aggregates the assigned tuples locally. However, this method suffers significantly from data skew. A simple solution is to use a local hash table to preaggregate heavy hitters before partitioning (Ye et al., 2011).
[…]
A sample execution of partitioned aggregation. This diagram is modified slightly from the one presented in (Leis et al., 2014) and is provided to contrast with the sample execution by our fully concurrent model on the same input in Figure 2.
Figure 1. A sample execution of partitioned aggregation. This diagram is modified slightly from the one presented in (Leis et al., 2014) and is provided to contrast with the sample execution by our fully concurrent model on the same input in Figure 2.
Leis et al. (Leis et al., 2014) use a partitioned aggregation strategy (originating in (Raman et al., 2013)) illustrated in Figure 1. This method consists of two stages: local preaggregation and partition-wise aggregation. First, each thread aggregates all values from their assigned morsel(s) in a fixed-sized local hash table ❶. When the hash tables are full, the partially aggregated values are spilled into partitions ❷ and the process continues. In the second stage, partitions are exchanged between threads, and each thread combines all the partial aggregates to compute the final answer ❸.
This algorithm is motivated by the need for skew resistance, avoiding uneven work distribution by spreading the work of aggregating heavy hitters among all threads in the local preaggregation stage. However, for high cardinality workloads, there is repeated spilling from the local aggregation tables. The constant spilling results in each tuple essentially being aggregated twice, once in each stage of the algorithm, creating a significant source of overhead.
The local preaggregation approach to partitioned aggregation has been adopted by several real-world systems, including DuckDB (Kuiper et al., 2024) and Datafusion (Lamb et al., 2023). Due to the widespread adoption of this particular algorithm and its good scaling behavior (Ye et al., 2011; Kuiper et al., 2024), it is used here as the baseline partitioned aggregation method.
@@ 2.3. Fully Concurrent Aggregation Model
[…]
A sample execution of our fully concurrent aggregation model using the same instance as in Figure 1. This particular diagram aligns most closely to our atomic or locking method of partial aggregate updates. In contrast to partitioned aggregation, instead of completing all work in local data structures, there are now two shared structures, the ticketing hash table and the vector of partial aggregates.
Figure 2. A sample execution of our fully concurrent aggregation model using the same instance as in Figure 1.
To perform a group by aggregation in a fully concurrent manner, each thread must aggregate an arbitrary morsel of data from start to finish, without any key distribution. We separate fully concurrent aggregation into two different steps: ticketing and partial aggregation update, as shown in Figure 2.
During ticketing ❶, a concurrent shared hash table is used to map each key value to an integer “ticket.” This mapping is one-to-one: each unique key is granted a single, unique ticket, and the ticket assigned to a key is consistent across all threads due to the use of a shared hash table. A ticket conceptually represents the location where the corresponding partial aggregate to be updated is stored, which in our implementation is simply an index of a vector. Alternatively, the ticketing step can be viewed as incrementally creating a perfect hash function. When we insert a new key, we also duplicate the key and store it in a vector in ticket order, which is later used to reconstruct the key column.
As a result of the first step, we have a morsel-sized column of tickets, stored as a vector. This column of tickets (T) and the input morsel of values (V) are fed to the update step ❷, where the query’s specified aggregation function(s) are applied (e.g., incremented for COUNT, added for SUM, etc). This step is the “aggregation” part of group by aggregation.
@@
Note that this layer of indirection that separates the partial aggregate from the hash map is already commonplace in practice. For example, it is used by MonetDB  (Boncz et al., 2005) to enable vectorized execution, as well as by DuckDB (Kuiper et al., 2024) and DataFusion (Lamb et al., 2023). Indirection also enables optimizations only possible when acting on a dense column of values, such as SIMD, or to amortize the cost of dynamic type resolution in non-compiled systems.
Given that indirection is already commonplace, we formalize it in our model to take advantage of its logical properties. In particular, since each unique key is assigned a ticket exactly once, our hash table workload is reduced to only lookups and inserts. We take advantage of this fact by designing a customized concurrent hash table that supports only these operations. This indirection also pushes concurrent updates to the second stage. Thus, concurrent updates act on a vector of values, rather than units of a hash table, allowing for alternate ways of synchronizing writes that do not degrade the performance of the ticketing hash table.
2.4. Experimental setup
Since our analysis and discussion will integrate several microexperiments, we first describe our experimental setup. Readers may wish to skip this section and reference it as needed.
We primarily run our experiments on a machine with 256 GB of RAM and an AMD EPYC 9454P processor with 48 cores @2.75GHz (“AMD”). This chip supports simultaneous multi-threading (SMT) up to 96 threads, has 3MB L1 cache, 48MB L2 cache, and a 256MB shared L3 cache. To test a diversity of architectures, we also run tests on an ARM system with an Ampere Altra Q80-30 @3.0GHz with 80 cores (no SMT support), 256GB RAM, 10MB L1 cache, and 80MB L2 cache (“ARM”), and an Intel system with an Intel Xeon Gold 5412U @2.10GHz with 24 cores (SMT up to 48 threads), 256GB RAM, 2MB L1 cache, 48MB L2 cache, 45MB shared L3 cache (“Intel”). For all platforms, clock boosting was disabled, as was address space layout randomization (which caused significant performance variation from run to run). All experiments were implemented in Rust and were compiled in release mode using the target-cpu=native flag.
Datasets
Our tests use synthetic datasets consisting of 100 million key-value pairs. Each key and value is a 64-bit integer (other than in Section 4.4). We varied the cardinality of the dataset, labeled low, high, and unique, which consist of 1000 unique keys (100% lookup/0% insertion on ticketing), 10 million unique keys (90% lookup/10% insertion), and 100 million unique keys (0% lookup/100% insertion) respectively, distributed perfectly uniformly. On the high cardinality dataset, we also add two types of skew: a Zipfian distribution with exponent parameter $s=0.8$, as well as a heavy-hitter dataset where 50% of the dataset consists of the same key.
Workload
As a demonstrative example of common aggregation functions, we use SUM for all tests. All experiments (except those in Section 4.5) assume perfect cardinality estimation and thus perfectly size hash tables and partial aggregate vectors. All results are obtained by taking the median latency of a given workload after 9 runs (not including warm-up runs).
@@ 3. Fully Concurrent Aggregation
Here, we investigate the design space for each stage of fully concurrent aggregation. We discuss methods for and execute micro-benchmarks of each stage in isolation to identify promising methods, which we later test end-to-end in Section 4.
@@ 3.1. Ticketing
When a new key arrives, an insert operation is required. Following a successful insertion of a key, all future requests must look up the inserted ticket value. In many concurrent hash table designs, this lookup operation can be achieved with much less overhead than an insert (e.g., shared locks are cheaper than exclusive locks). However, since we do not know ahead of time whether an insert is necessary (i.e., we do not know if a particular key has already been given a ticket), it is crucial to provide an efficient fast path for an atomic get-or-insert operation. Such an operation is rarely supported out-of-the-box by concurrent hash maps, but depending on the specific hash map implementation, the same effect could be achieved using an entry API or with a lookup followed by a non-overwriting insert.
@@
To efficiently retrieve keys in the same order as our aggregate values when materializing our final results, we also need the hash table to store a copy of keys in ticket order. This creates some performance overhead but is generally efficient since there is no contention–only the thread issuing a ticket is responsible for storing a copy. The ticket-order copy of keys can also be maintained as the only copy to decrease memory overhead (e.g., since the ticket value can be used to lookup the key during hash map operations), but would come at the cost of an additional cache miss per lookup.
@@
[…]
   (t,k):=(table[idx].t.load(),table.[idx].k.load())(t,k):=(table[idx].t.load(),table.[idx].k.load())
   if $t\geq 2$ then $\triangleright$ Fast path lookup.
@@
      return t2t-2
@@
   else if $table[idx].t.CAS(0,1)$ then $\triangleright$ Insert.
@@
     $table[idx].t.store(ticket+2)$
     return ticketticket
@@
Hash Table Designs
We benchmark a variety of state-of-the-art hash table designs. These include cuckoo hashing (Li et al., 2014), Iceberg hashing (Bender et al., 2023; Pandey et al., 2023), Rust’s Leapfrog (Clucas, [n.d.]) library, which uses leapfrog probing (Preshing, 2016), and Rust’s popular DashMap (Wejdenstål, [n.d.]).
We also implement a variant of the Folklore hash table proposed by Maier et al. (Maier et al., 2018), a lockless linear probing hash table. Our implementation leverages that we control the value being stored in the table (i.e., the ticket). We reserve 0 to indicate the key-value pair of the slot is empty and 1 to indicate a write is in progress. These properties allow our design to use only a single-word CAS instruction instead of the two-word version (which is not universally supported) required by the canonical implementation. We denote our implementation Folklore* to indicate the differences. The GET_OR_INSERT procedure for this design is outlined in Algorithm 1.
Our implementations additionally benefit from vectorized execution that optimizes hashing and amortizes the acquisition of a read lock on the shared table (which is needed to ensure the correctness of resizing).
[…]
Performance of the Folklore* hash table when using a fuzzy ticketer as opposed to an incrementing atomic counter to issue tickets. Performance is evaluated with both low and high cardinality workloads and is measured with latency (lower is better).
Figure 3. Performance of the Folklore* hash table with a fuzzy ticketer as opposed to an incrementing atomic counter. Performance is evaluated with latency (lower is better).
Micro Experiment: Generating Ticket Values
Generating a ticket value is simple on the surface, but actually presents a potential source of contention. Multiple threads must avoid issuing the same ticket to different keys. A naive implementation of a “ticketer” would use an atomic counter, incremented whenever a thread needs a new ticket. However, contention on the counter is high for insert-heavy workloads. To combat this issue, one can use a fuzzy ticketer that assigns each thread a range of tickets to issue at a time. Each thread only needs to concurrently access an atomic value when it has exhausted its assigned range. The tradeoff is that the vector of partial aggregates may no longer be perfectly dense, but the number of gaps is bounded linearly by the number of threads and does not cause much overhead to remove.
There is a large latency gap on insert-heavy workloads between using a pure atomic counter and our fuzzy ticketer, as seen in Figure 3 (lower latency is better). This micro-benchmark is tested on our best-performing hash table design, Folklore* at 48 threads. Latency in the high cardinality workload is greater for the pure atomic method by a factor of 9.0x, and even worse for unique keys. Since using a single atomic ticket value degrades performance when there are many unique keys, we recommend that implementers amortizes the cost of concurrent accesses across multiple inserts, such as with our fuzzy ticketer.
Micro Experiment: Memory Allocation
For memory-intensive database workloads such high-cardinality aggregations, the performance of the memory allocator can have a notable impact on the overall latency: a global hash table must be allocated. Prior work has compared various memory allocators based on their performance, memory efficiency, and scalability (Durner et al., 2019). The allocation of a large global hash table for high cardinalities is not only time-consuming, but is crucially also single-threaded. This single-threaded work is a constant cost that does not scale with threads, and can severely degrade performance at high thread counts. In Table 1, at 24 threads we can see that pre-optimization memory allocation can be more than half of total latency on some systems.
Table 1. Impact of zero-allocation on Folklore* performance on AMD and Intel at 24 threads with unique keys.
System AMD Intel
Zero-allocate No Yes No Yes
Total latency (sec) 0.48 0.38 0.92 0.27
Allocation (sec) 0.24 0.00 0.68 0.00
Ticketing (ec) 0.23 0.37 0.23 0.27
Zero-allocation was used to optimize the memory allocation step and improving the scaling of concurrent hash tables. Although zero-allocation is restricted to only certain types, this restriction is not a problem for the ticketing use case as long as the hash table is designed to not prematurely access uninitialized key fields. Zero-allocation, often a special type of call to the allocator (i.e., calloc in C), enables a number of optimizations by the operating system. A particularly important one is copy-on-write optimization, which defers allocation of a page to its first access rather than on the initial call to the allocator. As a result, memory allocation is no longer on the single-threaded critical path. When using zero-allocation, we see in Table 1 that the initial allocation becomes negligible in exchange for a modest performance hit during the actual ticketing step. Thus, we encourage practitioners to consider zeroed memory allocation strategies for large tables.
[…]